# Hole-spin qubits in germanium beyond the single-particle regime

### v1

@@ Introduction.
We hereafter investigate the effect of a homogeneous biaxial strain, such as the one that can result from the lattice mismatch in Ge/Si1-xGex heterostructures [Fig. 1(c)]. The Hamiltonian term related to a small and homogeneous strain is derived, within the $\bm{k}\cdot\bm{p}$ formalism, by the Bir-Pikus theory [77, 78, 79]. In particular, we consider a diagonal uniaxial strain tensor, $\varepsilon_{\beta}^{\alpha}=\delta_{\beta}^{\alpha}\varepsilon_{\alpha}^{\alpha}$, with $\varepsilon_{x}^{x}=\varepsilon_{y}^{y}=\varepsilon_{\parallel}$ and $\varepsilon_{z}^{z}=-2\varepsilon_{\parallel}C_{12}/C_{11}$, (where $C_{11}=126$ GPa and $C_{12}=44$ GPa are the elastic constants of Ge [80]), and a typical Ge concentration $x=0.8$ in the alloy, for which the measured strain parameter is $\varepsilon_{\parallel}=-0.0063$ [81].
Figure 3: (a) Dephasing time scale $\tau$ and (b) quality factor $Q=f_{{\rm R},x}\tau$ as functions of $\hbar\omega_{x}$, for $\theta=0^{\circ}$. The charge-noise induced electric field is oriented along the $z$ direction, with $\left|\bm{E}_{\rm cn}\right|=1$ mV///nm. The different curves correspond to: unstrained SHQ (red, solid); strained SHQ (red, dashed); unstrained THQ (blue, solid); strained THQ (blue, dashed).
In general, the presence of strain does not modify the qualitative picture outlined above. In fact, strain increases the energy difference between the HH and LH components, enhancing the HH character of the single-hole ground state. This results in an increased (decreased) $g$ factor for $\theta=0\degree$ ($\theta=90\degree$), both in the SHQ and in the THQ. As to the Rabi frequencies, their values are systematically reduced by the presence of strain, and to a similar extent for the SHQs and the THQs (Table 2). The peaks in the Rabi-frequency dependence on $\omega_{x}/\omega_{y}$ show up also in the strained system, albeit slightly displaced and narrowed with respect to the unstrained case (see the SM [74]).
Qubit dephasing.
Random fluctuations in the energy gap $\Delta_{10}$, induced by charge noise, are generally regarded as the main source of decoherence for hole-spin qubits [2]. The qubit sensitivity to electric fields can thus be quantified in terms of the dephasing time scale $\tau$, defined as:
$\frac{1}{\tau}\equiv\frac{e}{h}\Big{|}\bm{E}_{\rm cn}\cdot\left(\left\langle 1|\hat{\bm{r}}|1\right\rangle-\left\langle 0|\hat{\bm{r}}|0\right\rangle\right)\Big{|}\,.$ (4)
Here, $\bm{E}_{\rm cn}$ is a characteristic value of the charge-noise induced electric field vector. If the field is spatially homogeneous throughout the QD region then, due to the symmetries of the confinement potential, the dominant contribution in $1/\tau$ is the one related to the $z$ component of $\bm{E}_{\rm cn}$.
For the magnetic-field orientation $\theta=0^{\circ}$, which corresponds to the largest Rabi frequencies, the values of $\tau$ obtained for the THQ are smaller but comparable to those of the SHQ [Fig. 3(a)]. The biaxial strain approximately doubles the value of $\tau$ in both cases, thus increasing the difference between the two. In order to derive a more significant figure of merit, we introduce the quality factor $Q\equiv f_{{\rm R},x}\tau$. Its dependence on the QD in-plane aspect ratio, displayed in Fig. 3(b), leads to two main remarks. First, the increase in $\tau$ induced by the strain overcompensates the reduction in the Rabi frequency, resulting in an enhanced quality factor. Second, the THQ displays larger values of $Q$ in a wide range of values of $\hbar\omega_{x}$, especially as this becomes comparable to $\hbar\omega_{y}$. Overall, the relation between the quality factors of the SHQ and of the THQ confirms the suitability of a quasi-circular geometry for the implementation of the THQ.
@@ Conclusions.
The various scenarios considered in this work consistently point to the fact that three-hole spin qubits in Ge perform better than their single-hole counterparts when the QD is close to the circular regime. In both cases, the largest Rabi frequencies associated to in-plane oscillating electric fields are obtained for magnetic fields oriented along the growth direction. Biaxial strain leads to an overall reduction of the Rabi frequencies, which is however overcompensated by a proportional increase in the characteristic dephasing time, resulting in increased values of the quality factor.
@@ Acknowledgements.
We acknowledge the CINECA award under the ISCRA initiative, for the availability of high-performance computing resources and support (IsCb4-GERONIMO - HP10CG99PB). We thank the members of the Veldhorst, Niquet, Rimbach-Russ, and Scappucci groups for helpful discussions. We acknowledge financial support from NCCR Spin Grant No. 51NF40-180604 and from the PNRR Project PE0000023-NQSTI. SB was also supported by the EU through the H2024 QLSI2 project and the Army Research Office under Award Number: W911NF-23-1-0110. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the Army Research Office or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein.

### v2

@@ Introduction.
We hereafter investigate the effect of a homogeneous biaxial strain, such as the one that can result from the lattice mismatch in Ge/Si1-xGex heterostructures. The Hamiltonian term related to a small and homogeneous strain is derived, within the $\bm{k}\cdot\bm{p}$ formalism, by the Bir-Pikus theory [78, 79, 80]. In particular, we consider a diagonal uniaxial strain tensor, $\varepsilon_{\beta}^{\alpha}=\delta_{\beta}^{\alpha}\varepsilon_{\alpha}^{\alpha}$, with $\varepsilon_{x}^{x}=\varepsilon_{y}^{y}=\varepsilon_{\parallel}$ and $\varepsilon_{z}^{z}=-2\varepsilon_{\parallel}C_{12}/C_{11}$, (where $C_{11}=126$ GPa and $C_{12}=44$ GPa are the elastic constants of Ge [81]), and a typical Ge concentration $x=0.8$ in the alloy, for which the measured strain parameter is $\varepsilon_{\parallel}=-0.0063$ [82].
The presence of strain significantly affects the values of the quantities considered so far. The strain increases the energy difference between the HH and LH components, enhancing the HH character of the single-hole ground state. This results in an increased $g$-factor anisotropy, both in the SHQ and in the THQ. As to the Rabi frequencies, their values are generally reduced by the presence of strain, to a similar extent for the SHQs and the THQs. However, the presence of biaxial strain also results in a systematic increase of $\tau^{(1)}$ and $\tau^{(3)}$, as shown in Table 2 for the cases of the elongated and quasi-circular dots. Overall, the QD quality factor presents non-monotonic dependencies on the in-plane aspect ratio, with no general advantage or disadvantage resulting from the presence of strain (Fig. 5). On the other hand, the value of $Q_{x}$ obtained for the THQ largely exceeds that of the SHQ (figure inset) in the quasi-circular geometry, both for the strained and for the unstrained dot. This shows that the advantage of the three-hole encoding in terms of Rabi frequency ovecompensates the advantage of the SHQ at the level of dephasing time scale.
Figure 5: Quality factor $Q_{x}$ for the THQ as a function of $\hbar\omega_{x}$ (with $\hbar\omega_{y}=6\,$meV), for a magnetic field intensity B=0.05B=0.05\,T oriented out fo plane (parallel to the $z$ axis). Yellow and purple symbols correspond to the strained and to the unstrained dots, respectively. Inset: $Q_{x}$ vs $\hbar\omega_{x}$ for the SHQ, with the same units and the same color code as in the main plot.
@@ Conclusions.
The various scenarios considered in this work consistently point to the fact that three-hole spin qubits in Ge perform better than their single-hole counterparts when the QD is close to the circular regime. In both cases, the largest Rabi frequencies associated to in-plane oscillating electric fields are obtained for magnetic fields oriented along the growth direction. Biaxial strain leads to an overall reduction of the Rabi frequencies, which is however balanced by a comparable increase in the characteristic dephasing time, resulting in no significant difference between strained and unstrained dots in terms of quality factor.
@@ Acknowledgements.
We are particularly grateful to Y.-M. Niquet for an insightful comment, and we thank the members of the Niquet, Veldhorst, Rimbach-Russ, and Scappucci groups for helpful discussions. We acknowledge the CINECA award under the ISCRA initiative, for the availability of high-performance computing resources and support (IsCb4-GERONIMO - HP10CG99PB). This work was funded by the Italian PNRR Project PE0000023-NQSTI and, as a part of NCCR SPIN, a National Centre of Competence in Research, by the Swiss National Science Foundation (grant number 225153). SB was also supported by the EU through the H2024 QLSI2 project and the Army Research Office under Award Number: W911NF-23-1-0110. DL acknowledges the Deanship of Research at KFUPM and the Quantum Center for the support received under Grant no. CUP25102 and no. INQC2500, respectively. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the Army Research Office or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein.